# Physical properties of circumnuclear ionising clusters. III. Kinematics of gas and stars in NGC 7742

### v2

@@ 4.2 Dynamical masses
We have calculated the dynamical masses for each observed cluster from its measured CaT velocity dispersion and size assuming a virialised system and using the expression given by (Ho & Filippenko, 1996a, b):
where is the one-dimensional velocity dispersion of the system, R is its radius and G is the gravitational constant.
This calculation assumes that the clusters (i), have spherical symmetry; (ii), are gravitationally bound; and (iii), have an isotropic velocity distribution. Most of the clusters show only one knot at thus the calculated masses under this assumption seem to be appropriate. However a small number are clearly multiple, in this case the total mass of the cluster was estimated by adding the masses of its individual knots inferred using the velocity dispersion of the entire region a procedure that most likely result on an overestimate.
Figure 10 shows the distribution of the cluster dynamical masses for NGC 7742 compared with those of the inner and outer rings of NGC 7469. We can see that NGC 7742 knots show lower masses which is to be expected from their smaller sizes. However, they tend to be more massive than CNSFRs from the literature (Hägele et al., 2010, 2007, 2009, NGC 3351, NGC 2903, and NGC 3310).
@@
Table 7 lists for each region in columns 1 to 5: (1) the region ID; (2) the number of stellar knots present in the region, (3) the mean radius in pc, (4) the dynamical mass in M⊙ and (5) the ratio of ionising to dynamical mass. To calculate this ratio, we have used the ionising masses calculated in the first paper of the series from the number of hydrogen ionising photons, Q(H), and the equivalent width of the H emission line to take into account the cluster evolution. We used a Salpeter IMF , with , = 0.85 and = 120 that seems the most suitable for the age of our regions.
Figure 11 shows the distribution of the calculated dynamical masses for each CNSFR. The dynamical masses of inner and outer ring regions of NGC 7469, calculated in the second paper of this series are also shown. Additionally, we show the masses of CNSFRs reported in the literature for NGC 3351, NGC 2903, and NGC 3310, which have been calculated with the same method described in this work (Hägele et al., 2010, 2007, 2009). The dynamical masses of NGC 7742 clusters range from 2.5 106 to 1.0 108 M⊙, values similar to those found in the rest of the galaxies shown but lower than those obtained for the two rings of NGC 7469.
The ratio of ionising to dynamical masses has values ranging from 0.15 to 7.07 % implying a contribution by recent star formation for the ring regions larger than found in other galaxies. For comparison, this percentage ranges from 0.02 to 0.91 % for the CNSFRs in the outer ring of NGC 7469 whose distance to the galactic nucleus is similar to that of the NGC 7742 ring to its nucleus.
@@ 4.3 Kinematics of stars and ionised gas
In Section 3.2 we have derived the radial velocities and velocity dispersions of the ionised gas in the apertures of the selected HII regions. Our results suggest the existence of at least two kinematically distinct H components that can be identified in the spectra of CNSFR with high S/N. We have compared their radial velocities in order to find out which of them are associated with the ionising clusters.
@@
The left panel of Figure 12 shows the comparison of the velocities of the H broad and narrow components with the 1 to 1 line represented by a dashed grey line. We can see that both radial velocities coincide within the errors. Only two regions seem to fall off the 1 to 1 line: R62, and R65. An additional component of the H emission line has been identified in 7 of the regions and a comparison of their velocities and those of the narrow component is shown in the right panel of the same figure. Only three of these additional components, in regions R1, R7, and R67, seem to be associated with the ionising clusters, hence they can be attributed to local effects inside the regions. In the other four cases that component may have an external origin.
@@
The left panels of Figure 13 show the gas velocity of the narrow component of forbidden emission lines as a function of the corresponding H one. Results for the [NII]6584 Å, [SII]6731 Å, and [SIII]9532 Å emission lines are shown from top to bottom. All regions follow the 1 to 1 relation within the errors, hence assuring that the selected kinematical components are associated to the observed HII regions. Only in the case of the [SIII] emission line a few regions seem to deviate from the identity relation; these regions show large observational errors related to low S/N spectra. The right panels of the figure show the same relations but for the broad line component of both forbidden and H lines. All [NII] broad components follow the 1 to 1 relation, their association with the observed HII regions is guaranteed; for the [SII] lines, only in three of the regions the presence of a broad component is consistent with the S/N ratio, and only one region shows a broad component in the [SIII] emission lines; this region, R1, displays a complex kinematics with three different width velocity components, clearly identified in the H and [NII]6548,84 Å lines, that seem to be associated with the HII region (see Figure 3).
@@
Figure 14 shows the comparison between the derived velocity dispersion for stars and gas in the selected CNSFRs. Gas narrow and broad components are represented by blue dots and green triangles respectively while regions showing a single component are represented by purple squares. The dashed grey line shows the 1 to 1 relation. The upper left panel of this figure shows the results for the H emission line. We can see that the narrow components have velocity dispersions lower than the stellar ones and in a narrow range with a mean value of = 9.26 km/s, shown by a red solid line with dotted red lines representing the standard deviation of the distribution. Regions which have only one component show a similar behaviour. On the other hand, the broad components lie slightly above the 1 to 1 relation in regions with low star velocity dispersions. This behaviour seems to be lost for higher values of . The upper right panel of the figure shows the results for [NII] having the same behaviour as the H line. In this case, the mean velocity dispersion of the narrow component is = 11.2 km/s and the standard deviation is slightly larger. For the broad component the same trend seen for the H line is easily recognizable. The only component identified in the [SII] lines, shows the same velocity dispersion associated to the gas narrow component, showing a mean value of = 15.8 km/s (see the lower right panel of the figure) while for the [SIII] lines velocity dispersions show a more complex behaviour as shown in the lower right panel of the figure; the single component present shows a wider velocity dispersion distribution, although mostly below the 1 to 1 line. The only region with a broad component in [SIII] is R1 appearing in the upper left corner of the diagram, with a gas velocity dispersion of 65 km/s (this region is close to supernova SN 1993R).
@@
Similar results to those presented above have also been found in other extragalactic HII region studies in the literature. For example in Firpo et al. (2010), where some giant HII regions in the NGC 7479 spiral galaxy were analysed using echelle spectroscopy data with a similar spectral resolution to that of MEGARA. In this work, region III is reported to show a broad Gaussian component in the H and [NII] emission lines with a velocity dispersion in the range between 37.8 km/s and 51.2 km/s while the narrow one is around 10 km/s and no broad component was found in the [OIII] profile whose velocity profile can be fitted with a Gaussian with a value of around 20 km/s. Also, CNSFRs of different galaxies studied in Hägele et al. (2009, 2010) show broad components in the emission lines at the same radial velocity as the narrow ones within the errors; the authors identified the narrow and broad components with ionised gas supported by rotation and dynamical pressure respectively. Interestingly, the CNSFR rings in both galaxies are thought to have been produced in a minor merger event.
@@
A diversity of line broadening mechanisms have been proposed to explain the line profiles of HII regions. Winds associated with WR stars or young main sequence O stars (i. e. Diaz et al., 1987; Terlevich et al., 1996) could produce them. However, the expected broadening due to WR winds would, in general, be larger than observed in our clusters since these stars can have winds of thousands of km/s. On the other hand, we have estimated the number of O stars that may be present in the CNSFRs of this study. For that, we have used the Salpeter (1955, mlow = 0.85M⊙, mup = 120 M⊙) IMF and models from (Smith et al., 2002, Z⊙, log(g) = 4, R∗ = 7.1 - 10.6 R⊙) deriving the following equation:
@@
where Mcluster is expressed in solar masses. For the clusters with derived dynamical mass this number is between 1 and 38, with a mean value of 10 stars. However, normal O stars have winds with velocities 103 km/s, depending on the density of the environment. Thus, they would also provide a broadening larger than observed. In conclusion, we rule out stellar winds as the main broadening mechanism in our CNSFRs.
@@
Other possible explanation proposed in the literature is the turbulence of the interstellar gas. In 1994, Chu & Kennicutt (1994) studied the 30 Doradus giant HII region, in the Large Magellanic Cloud. The outer regions of the nebula have a smooth turbulent velocity field with velocities of 13 to 17 km/s and, at the central regions, the velocity is dominated by multiple expanding structures from 8.5 to 85 km/s (in units of ). They corroborated that the integration of all these structures produced a simple broad Gaussian profile with faint extended wings. Additionally, several fast expanding shells were found with velocities from 40 to 130 km/s. (Medina Tanco et al., 1997) studied the giant HII region NGC 604, in M33 finding structure in the velocity field projected along the line of sight that could be attributed to turbulence and suggesting that this turbulence could be powered by recent star formation processes. According to this, the broad component of velocity dispersions found in our particular case look similar to those produced by interstellar gas turbulence.
However, the works just mentioned above lack the information related to the comparison between gas and stars velocity dispersions and this fact provides additional important information: the stars and the gas associated to the broad component of the emission lines showing similar values of velocity dispersion, sharing a common kinematics, can be explained if they respond to the same gravitational potential of the cluster. From the results obtained in the first paper of this series, we concluded that ring clusters are composed by a young non-ionising intermediate age population about 300 Ma old, plus an ionising population of 4.7 Ma in a 4:1 ratio. In this scenario, the ring clusters would have formed almost simultaneously in a first star formation episode. The stars in such a population would return part of their gas to the interstellar medium during their evolution. A fraction of this gas could escape from the cluster, while some of it would be unable to do so thus remaining trapped in the cluster (Silich & Tenorio-Tagle, 2018). This gas would feel the same gravitational potential as the cluster stars, consistent with our results. A fraction of the gas that cooled down would form a new stellar population, which would be responsible for the ionising photons. In this scenario, the narrow gas component would be associated with the HII region ionisation shell. According to the results presented here (see Sec. 4.2, this young ionising stellar population accounts for between 0.15 and 7.07 % of the total dynamical mass of the cluster.
@@ 5 Conclusions
In this third paper of a series, we analyse the kinematics of gas and stars in the circumnuclear ring of NGC 7742 estimating the dynamical masses of their CNSFRs from high spectral resolution data.
We have extracted spectra for each region using the apertures selected from H images.
The analysis of the emission line profiles, show the presence of two different kinematical components with the same radial velocity. These two different components have been found in the H emission line in 90 % of the analysed regions with most of them showing the same component also in the [NII] emission lines. The narrow component has an almost constant value between 10 km/s and 15 km/s while the broad component velocities are between 20 and 45 km/s.
@@
The sizes of the star clusters have been measured from a F555W WFPC2-HST image. About 60 % show a single star forming knot with the rest of them showing between 2 and 4 knots each. Cluster velocity dispersions have been derived from the CaT absorption lines using a modified version of Tonry & Davis’ cross-correlation technique. These velocities together with the measured sizes have been used to calculate the cluster dynamical masses under the assumption that the clusters are virialised. The masses have been found to be between 2.5 106 and 1.0 108 M⊙, values similar to those found in the literature for circumnuclear clusters.
Finally, a comparison of the gas and star velocity dispersion, seldom found in the literature, has been made revealing a complex internal structure of the gas. A scenario has been put forward suggesting that the circumnuclear ring clusters would have formed simultaneously in a first star formation episode whose high mass stars would have returned part of their gas to the interstellar medium during their evolution. A fraction of this initial ejecta remains trapped in the gravitational potential of the cluster. A fraction of the gas trapped would cooled down and form the young stellar population responsible for the ionisation of the gas currently observed. This young stellar population accounts for between 0.15 and 7.07 % of the total dynamical mass of the cluster.
@@ Acknowledgements.
This research has made use of the services of the ESO Science Archive Facility and NASA’s Astrophysics Data System Abstract Service. It is based on observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programme 60.A-9301(A) and data products created thereof. Also we have used observations obtained with the NASA/ESA HST and also from the Hubble Legacy Archive, which is a collaboration between the Space Telescope Science Institute (STScI/NASA), the Space Telescope European Coordinating Facility (ST-ECF/ESA), and the Canadian Astronomy Data Centre (CADC/NRC/CSA). This work has been supported by Spanish grants from the former Ministry of Economy, Industry and Competitiveness through the MINECO-FEDER research grant AYA2016-79724-C4-1-P, and the present Ministry of Science and Innovation through the research grant PID2019-107408GB-C42. S.Z. acknowledges the support from contract: BES-2017-080509 associated to the first of these grants. This work is based on observations made with the Gran Telescopio Canarias (GTC), installed at the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias, on the island of La Palma. The data was obtained with the instrument MEGARA funded by European Regional Development Funds (ERDF), through Programa Operativo Canarias FEDER 2014–2020.References

### v3

@@ 4.2 Dynamical masses
We calculated the dynamical masses for each observed cluster from its measured CaT velocity dispersion and size assuming a virialised system and using the expression given by (Ho & Filippenko, 1996a, b)
where is the one-dimensional velocity dispersion of the system, R is its radius, and G is the gravitational constant.
This calculation assumes that the clusters (i) have spherical symmetry; (ii) are gravitationally bound; and (iii) have an isotropic velocity distribution. Most of the clusters show only one knot, and thus the calculated masses under this assumption seem to be appropriate. However, a small number are clearly multiple; in this case the total mass of the cluster was estimated by adding the masses of its individual knots inferred using the velocity dispersion of the entire region a procedure that most likely result in an overestimate.
Figure 10 shows the distribution of the cluster dynamical masses for NGC 7742 compared with those of the inner and outer rings of NGC 7469. We can see that NGC 7742 knots show lower masses, which is to be expected from their smaller sizes. However, they tend to be more massive than the CNSFRs from the literature (Hägele et al., 2010, 2007, 2009, NGC 3351, NGC 2903, and NGC 3310).
@@
Table 7 lists for each region in columns 1 to 5: (1) the region ID; (2) the number of stellar knots present in the region; (3) the mean radius in pc; (4) the dynamical mass in M⊙; and (5) the ratio of ionising to dynamical mass. To calculate this ratio, we used the ionising masses calculated in the first paper of the series from the number of hydrogen ionising photons, Q(H), and the equivalent width of the H emission line to take into account the cluster evolution. We used a Salpeter IMF , with , = 0.85, and = 120, which seems the most suitable for the age of our regions.
Figure 11 shows the distribution of the calculated dynamical masses for each CNSFR. The dynamical masses of inner and outer ring regions of NGC 7469, calculated in the second paper of this series are also shown. Additionally, we show the masses of CNSFRs reported in the literature for NGC 3351, NGC 2903, and NGC 3310, which were calculated with the same method described in this work (Hägele et al., 2010, 2007, 2009). The dynamical masses of NGC 7742 clusters range from 2.5 106 to 1.0 108 M⊙, values similar to those found in the rest of the galaxies shown, but lower than those obtained for the two rings of NGC 7469.
The ratio of ionising to dynamical masses has values ranging from 0.15 to 7.07 %, implying a contribution by recent star formation for the ring regions larger than found in other galaxies. For comparison, this percentage ranges from 0.02 to 0.91 % for the CNSFRs in the outer ring of NGC 7469 whose distance to the galactic nucleus is similar to that of the NGC 7742 ring to its nucleus.
@@ 4.3 Kinematics of stars and ionised gas
In Sect. 3.2 we derive the radial velocities and velocity dispersions of the ionised gas in the apertures of the selected HII regions. Our results suggest the existence of at least two kinematically distinct H components that can be identified in the spectra of CNSFRs with high S/N. We compared their radial velocities in order to find out which of them are associated with the ionising clusters.
@@
The left panel of Fig. 12 shows the comparison of the velocities of the H broad and narrow components with the one-to-one line represented by a dashed grey line. We can see that the two radial velocities coincide within the errors. Only two regions seem to fall off the one-to-one line: R62, and R65. An additional component of the H emission line has been identified in seven of the regions, and a comparison of their velocities and those of the narrow component is shown in the right panel of the same figure. Only three of these additional components, in regions R1, R7, and R67, seem to be associated with the ionising clusters, and hence they can be attributed to local effects inside the regions. In the other four cases, that component may have an external origin.
@@
The left panels of Fig. 13 show the gas velocity of the narrow component of forbidden emission lines as a function of the corresponding H component. Results for the [NII]6584 Å, [SII]6731 Å, and [SIII]9532 Å emission lines are shown from top to bottom. All regions follow the one-to-one relation within the errors, hence assuring that the selected kinematical components are associated with the observed HII regions. Only in the case of the [SIII] emission line do a few regions seem to deviate from the identity relation; these regions show large observational errors related to low S/N spectra. The right panels of the figure show the same relations, but for the broad-line component of both forbidden and H lines. All the [NII] broad components follow the one-to-one relation; their association with the observed HII regions is guaranteed. For the [SII] lines, only in three of the regions is the presence of a broad component consistent with the S/N, and only one region shows a broad component in the [SIII] emission lines; this region, R1, displays complex kinematics with three different width velocity components, clearly identified in the H and [NII]6548,84 Å lines, that seem to be associated with the HII region (see Fig. 3).
@@
Figure 14 shows the comparison between the derived velocity dispersion for stars and gas in the selected CNSFRs. Gas narrow and broad components are represented by blue dots and green triangles, respectively, while regions showing a single component are represented by purple squares. The dashed grey line shows the one-to-one relation. The top left panel of this figure shows the results for the H emission line. We can see that the narrow components have velocity dispersions lower than the stellar values and in a narrow range with a mean value of = 9.26 km/s, shown by a red solid line; the dotted red lines represent the standard deviation of the distribution. Regions that have only one component show a similar behaviour. On the other hand, the broad components lie slightly above the one-to-one relation in regions with low star velocity dispersions. This behaviour seems to be lost for higher values of . The top right panel of the figure shows the results for [NII], which show the same behaviour as the H line. In this case, the mean velocity dispersion of the narrow component is = 11.2 km/s and the standard deviation is slightly larger. For the broad component the same trend seen for the H line is easily recognizable. The only component identified in the [SII] lines, shows the same velocity dispersion associated with the gas narrow component, showing a mean value of = 15.8 km/s (see the lower right panel of the figure), while for the [SIII] lines the velocity dispersions show a more complex behaviour, as shown in the lower right panel of the figure; the single component present shows a wider velocity dispersion distribution, although mostly below the one-to-one line. The only region with a broad component in [SIII] is R1 appearing in the top left corner of the diagram, with a gas velocity dispersion of 65 km/s (this region is close to supernova SN 1993R).
@@
Similar results to those presented above have also been found in other extragalactic HII region studies in the literature. For example in Firpo et al. (2010), where some giant HII regions in the NGC 7479 spiral galaxy were analysed using echelle spectroscopy data with a similar spectral resolution to that of MEGARA. In this work, region III is reported to show a broad Gaussian component in the H and [NII] emission lines with a velocity dispersion in the range between 37.8 km/s and 51.2 km/s, while the narrow component is around 10 km/s and no broad component was found in the [OIII] profile whose velocity profile can be fitted with a Gaussian with a value of around 20 km/s. In addition, CNSFRs of different galaxies studied in Hägele et al. (2009, 2010) show broad components in the emission lines at the same radial velocity as the narrow ones within the errors; the authors identified the narrow and broad components with ionised gas supported by rotation and dynamical pressure respectively. Interestingly, the CNSFR rings in both galaxies are thought to have been produced in a minor merger event.
@@
A diversity of line broadening mechanisms have been proposed to explain the line profiles of HII regions. Winds associated with WR stars or young main sequence O stars (i. e. Diaz et al., 1987; Terlevich et al., 1996) could produce them. However, the expected broadening due to WR winds would, in general, be greater than observed in our clusters since these stars can have winds of thousands of km/s. On the other hand, we estimated the number of O stars that may be present in the CNSFRs of this study. To do so we used the Salpeter (1955, mlow = 0.85M⊙, mup = 120 M⊙) IMF and models from (Smith et al., 2002, Z⊙, log(g) = 4, R∗ = 7.1 - 10.6 R⊙), which derive the following equation:
@@
Here Mcluster is expressed in solar masses. For the clusters with derived dynamical mass this number is between 1 and 38, with a mean value of 10 stars. However, normal O stars have winds with velocities 103 km/s, depending on the density of the environment. Thus, they would also provide a broadening larger than observed. In conclusion, we rule out stellar winds as the main broadening mechanism in our CNSFRs.
@@
Another possible explanation proposed in the literature is the turbulence of the interstellar gas. In 1994, Chu & Kennicutt (1994) studied the 30 Doradus giant HII region, in the Large Magellanic Cloud. The outer regions of the nebula have a smooth turbulent velocity field with velocities of 13 to 17 km/s and, at the central regions, the velocity is dominated by multiple expanding structures from 8.5 to 85 km/s (in units of ). They corroborated that the integration of all these structures produced a simple broad Gaussian profile with faint extended wings. Additionally, several fast expanding shells were found with velocities from 40 to 130 km/s. (Medina Tanco et al., 1997) studied the giant HII region NGC 604, in M33 finding structure in the velocity field projected along the line of sight that could be attributed to turbulence and suggesting that this turbulence could be powered by recent star formation processes. According to this, the broad component of velocity dispersions found in our particular case look similar to those produced by interstellar gas turbulence.
However, the works just mentioned above lack the information related to the comparison between gas and stars velocity dispersions, and this fact provides additional important information: the stars and the gas associated with the broad component of the emission lines showing similar values of velocity dispersion, and sharing a common kinematics, can be explained if they respond to the same gravitational potential of the cluster. From the results obtained in the first paper in this series, we concluded that ring clusters are composed of a young non-ionising intermediate age population about 300 Ma old, plus an ionising population of 4.7 Ma in a 4:1 ratio. In this scenario, the ring clusters would have formed almost simultaneously in a first star formation episode. The stars in such a population would return part of their gas to the interstellar medium during their evolution. A fraction of this gas could escape from the cluster, while some of it would be unable to do so thus remaining trapped in the cluster (Silich & Tenorio-Tagle, 2018). This gas would feel the same gravitational potential as the cluster stars, consistent with our results. A fraction of the gas that cooled down would form a new stellar population, which would be responsible for the ionising photons. In this scenario, the narrow gas component would be associated with the HII region ionisation shell. According to the results presented here (see Sect. 4.2, this young ionising stellar population accounts for between 0.15 and 7.07 % of the total dynamical mass of the cluster.
@@ 5 Conclusions
In this third paper in the series, we analysed the kinematics of gas and stars in the circumnuclear ring of NGC 7742 estimating the dynamical masses of their CNSFRs from high spectral resolution data. We extracted spectra for each region using the apertures selected from H images. The analysis of the emission line profiles, show the presence of two different kinematical components with the same radial velocity. These two different components were found in the H emission line in 90 % of the analysed regions; most of them also showed the same component in the [NII] emission lines. The narrow component has an almost constant value between 10 km/s and 15 km/s, while the broad component velocities are between 20 and 45 km/s.
@@
The sizes of the star clusters were measured from a F555W WFPC2-HST image. About 60 % show a single star-forming knot; the rest of them show between two and four knots each. Cluster velocity dispersions were derived from the CaT absorption lines using a modified version of the Tonry & Davis cross-correlation technique. These velocities together with the measured sizes were used to calculate the cluster dynamical masses under the assumption that the clusters are virialised. The masses were found to be between 2.5 106 and 1.0 108 M⊙, values similar to those found in the literature for circumnuclear clusters.
Finally, a comparison of the gas and star velocity dispersion, seldom found in the literature, was made revealing a complex internal structure of the gas. A scenario was put forward suggesting that the circumnuclear ring clusters formed simultaneously in a first star formation episode whose high-mass stars could have returned part of their gas to the interstellar medium during their evolution. A fraction of this initial ejecta remained trapped in the gravitational potential of the cluster. A fraction of the trapped gas cooled down and formed the young stellar population responsible for the ionisation of the gas currently observed. This young stellar population accounts for between 0.15 and 7.07 % of the total dynamical mass of the cluster.
@@ Acknowledgements.
This research has made use of the services of the ESO Science Archive Facility and NASA’s Astrophysics Data System Abstract Service. It is based on observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programme 60.A-9301(A) and data products created thereof. Also we have used observations obtained with the NASA/ESA HST and also from the Hubble Legacy Archive, which is a collaboration between the Space Telescope Science Institute (STScI/NASA), the Space Telescope European Coordinating Facility (ST-ECF/ESA), and the Canadian Astronomy Data Centre (CADC/NRC/CSA). This work has been supported by Spanish grants from the former Ministry of Economy, Industry and Competitiveness through the MINECO-FEDER research grants AYA2016-79724-C4-1-P, and PID2019-107408GB-C42 and the present Ministry of Science and Innovation through the research grant PID2022-136598NB-C33 by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. S.Z. acknowledges the support from contract: BES-2017-080509 associated with the first of these grants. This work is based on observations made with the Gran Telescopio Canarias (GTC), installed at the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias, on the island of La Palma. The data was obtained with the instrument MEGARA funded by European Regional Development Funds (ERDF), through Programa Operativo Canarias FEDER 2014–2020.References
- Böker et al. (2008) Böker, T., Falcón-Barroso, J., Schinnerer, E., Knapen, J. H., & Ryder, S. 2008, AJ, 135, 479
- Brandl et al. (2012) Brandl, B. R., Martín-Hernández, N. L., Schaerer, D., Rosenberg, M., & van der Werf, P. P. 2012, A&A, 543, A61
@@
- Knapen (2005) Knapen, J. H. 2005, A&A, 429, 141
- Levy et al. (2022) Levy, R. C., Bolatto, A. D., Leroy, A. K., et al. 2022, ApJ, 935, 19
- Lopez et al. (2023) Lopez, S., Lopez, L. A., Nguyen, D. D., et al. 2023, ApJ, 942, 108
- Mazzuca et al. (2008) Mazzuca, L. M., Knapen, J. H., Veilleux, S., & Regan, M. W. 2008, ApJS, 174, 337
@@
- Sarzi et al. (2007) Sarzi, M., Allard, E. L., Knapen, J. H., & Mazzuca, L. M. 2007, MNRAS, 380, 949